Transverse waves travel at 20.0 m/s on a string that is under a tension of 6.00 N. What tension is required for a wave speed of 30.0m/s in the same string
Answer #1
This is your first qn. Do rate the answers you get and keepyour respect % at 100%.
The relation is,
               velocity, v = √(T/m)
Linear density, m is the same.
Squaring and dividing,
                                  (v1/v2)2 =T1/T2
Tension,
         T2= T1((v2/v1)2 =6.0x(30/20)2
              = 6.0x9/4 = 13.5N